OpenStudy (anonymous):

Consider the following set of equations: Equation M: y = x + 4 Equation P: y = 3x + 6 Which of the following is a step that can be used to find the solution to the set of equations? A. x = 3x +6 B. x + 4 = 3x + 6 C. x + 6 = 3x + 4 D. x = 3x

pooja195 (pooja195):

Hi Welcome To Openstudy ^.^ So basically we are given 2 equations using the substitution method we can plug equation 1 into equation 2 or revers doesn't really matter what do we get once we put equation 1 into equation 2? y = x + 4 y = 3x + 6

OpenStudy (anonymous):

3x and x ?

pooja195 (pooja195):

Now take the first equation x+4 and replace the 2nd equation with x+4 instead of y what do we get?

OpenStudy (anonymous):

x+4=3x + 6 ?

pooja195 (pooja195):

Well done =)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)